The millers want to buy a mobile home that cost 36000.
They have already saved 27000. What percen of the total cost
have they saved.

Answer by jim_thompson5910(35256)

First divide: 27000/36000 = 27/36 = 3/4 = 0.75


Then multiply by 100 to convert to a percentage: 100*0.75 = 75%


They have saved 75% of the total cost so far.